Status Report

NASA STS-120 Execute Package FD 07

By SpaceRef Editor
November 1, 2007
Filed under , , ,
NASA STS-120 Execute Package FD 07

Full report

1. After maneuvering the SRMS to pre-cradle on FD6 and deselecting RMS, the shoulder 4 temps immediately started decreasing. Combined with the current ISS attitude, the temps are now acceptable. We are back to the nominal SRMS plan and request that RMS SELECT – OFF should be performed anytime the SRMS is at the pre-cradle position to keep the temps as low as possible.

2. Shuttle Condensate Collection Details – Shuttle Crew At MET 5/16:00, perform CHANGEOUT of SHUTTLE CONDENSATE COLLECTION (ORB OPS, 21 ECLS) p. 5-36 using empty CWC s/n 5057 which can be found in MF71M. CWC s/n 5104 can be temp stowed as it will be dumped on FD9. Unused DIDB Stowage

MCC would like for you to stow the two DIDB’s filled with hot water on FD3 in MF14H after sealing each bag in a Ziploc. Please stow the bags in any available MF14H tray not containing the saliva and blood collection return samples.

3. Request for Scott concerning EVA 2 DCS 760 Camera without Flash,

Photo/TV is continuing to troubleshoot the EVA Camera issues from EVA 2. Please clarify if you had any positive indications that your DCS 760 camera (Camera without Flash) was taking pictures (i.e. Shutter Release, Mirror Flip Up or Light-Pipe Indication).

4. Last night we power cycled the DAIU and reset ICOM A from the ground. When time permits, please perform another comm check on ICOM A and advise MCC of the results.

SpaceRef staff editor.